Three different machines A1, A2, A3 were used for producing a large batch of similar manufactured items. Suppose that 10% of items are produced by machine A1, 20% by A2, and 40 % by A3. Suppose further that 1% of the items produced by A1, 2% by A2, and 3% by A3 machine are defective. Suppose one item is selected from the entire batch and it is found to be defective. Calculate the probability that this item was produced by machine A2.


1
Expert's answer
2021-05-02T08:27:53-0400

Three different machines "A_1,A_2,A_3" were used for producing a large batch of similar manufactured items. Suppose that 40% of items are produced by machine "A_1," 20% by "A_2," and 40 % by "A_3." Suppose further that 1% of the items produced by "A_1," 2% by "A_2," and 3% by "A_3" machine are defective. Suppose one item is selected from the entire batch and it is found to be defective. Calculate the probability that this item was produced by machine "A_2."


Let "D=\{item\ is \ defective\}"

"A_1=\{item\ is \ produces\ by\ machine\ A_1\}"

"A_2=\{item\ is \ produces\ by\ machine\ A_2\}"

"A_3=\{item\ is \ produces\ by\ machine\ A_3\}"

We know that "P(A_1)=0.4,P(A_2)=0.2," and "P(A_3)=0.4."

Also

"P(D|A_1)=0.01,P(D|A_2)=0.02, P(D|A_3)=0.02"

By the Bayes’ theorem


"P(A_2|D)"

"=\dfrac{P(A_2)P(D|A_2)}{P(A_1)P(D|A_1)+P(A_2)P(D|A_2)+P(A_3)P(D|A_3)}"


"=\dfrac{0.2(0.02)}{0.4(0.01)+0.2(0.02)+0.4(0.03)}=0.2"

The probability that this item was produced by machine "A_2" is "0.2."